OpenStudy (anonymous):

\[In \triangle ABC(A2−B2)/(A2+B2)=\sin(A−B)/\sin(A+B)then the △ is\]

OpenStudy (dan815):

is it exactly like that no brackets missing anywhere right

ganeshie8 (ganeshie8):

(a^2-b^2)/(a^2\(\color{Red}{+}\)b^2) =sin(a-b)/sin(a+b) like this ?

OpenStudy (dan815):

ooo

OpenStudy (dan815):

is he gone xD

ganeshie8 (ganeshie8):

@dg2

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

can u edit the quesiton and fix typo ?

ganeshie8 (ganeshie8):

also use parenthesis so that it is less ambiguous

ganeshie8 (ganeshie8):

also use capital letters for angles : A,B,C and small letters for sides : a,b,c

ganeshie8 (ganeshie8):

I have modified the question, see if it looks okay now

OpenStudy (anonymous):

\[In \triangle ABC (A^2-B^2)/(A^2+B^2)=\sin(A-B)/\sin(A+B) then the \triangle is\]

OpenStudy (anonymous):

now okay .

ganeshie8 (ganeshie8):

copy paste below : ``` In \(\triangle ABC\) , \[ \dfrac{a^2-b^2}{a^2+b^2} = \dfrac{\sin(A-B)}{\sin(A+B)} \] then the \(\triangle \) is ? ```

OpenStudy (anonymous):

\[In \(\triangle ABC\) , \[ \dfrac{a^2-b^2}{a^2+b^2} = \dfrac{\sin(A-B)}{\sin(A+B)} \] then the \(\triangle \) is ?\]

ganeshie8 (ganeshie8):

don't use any brackets, copy past exact same stuff

OpenStudy (anonymous):

In \(\triangle ABC\) , \[ \dfrac{a^2-b^2}{a^2+b^2} = \dfrac{\sin(A-B)}{\sin(A+B)} \] then the \(\triangle \) is ?

ganeshie8 (ganeshie8):

finally ! congrats !!

ganeshie8 (ganeshie8):

familiar with `componendo and dividendo` ?

OpenStudy (anonymous):

yes .i got struck with this site .(a^2-b^2)/(a^+b^2)*(a^+b^2)/(a^+b^2)

OpenStudy (anonymous):

oh sorry sin(A-B)/sin(A+B)*sin(A+B)/sin(A+B)

ganeshie8 (ganeshie8):

componendo and dividendo : If \[\dfrac{p}{q} = \dfrac{r}{s}\] then : \[\dfrac{p+q}{p-q} = \dfrac{r+s}{r-s}\]

ganeshie8 (ganeshie8):

Since you're given : \[\dfrac{a^2-b^2}{a^2+b^2} = \dfrac{\sin(A-B)}{\sin(A+B)}\] applying componendo and dividendo, you get : \[\dfrac{(a^2-b^2) + (a^2 + b^2)}{(a^2-b^2) + (a^2 + b^2) } = \dfrac{\sin(A-B) + \sin(A+B)}{\sin(A-B) - \sin(A+B)}\]

OpenStudy (anonymous):

latex is not working for me

ganeshie8 (ganeshie8):

simplifying a bit, you would get : \[\large \dfrac{a^2}{b^2 } = \dfrac{\sin A \cos B}{\cos A \sin B}\]

ganeshie8 (ganeshie8):

here is a screenshot http://www.webpagescreenshot.info/img/53f70852d9c8f7-85003353

OpenStudy (anonymous):

screenshot also not loaded;;;) i understood a^2/b^2=(sina*cosb)/(cosa*sinb)

OpenStudy (anonymous):

so tan a*cot b ?

ganeshie8 (ganeshie8):

yes a^2/b^2 = tanA cotB

ganeshie8 (ganeshie8):

what to conclude from here ? hmm

ganeshie8 (ganeshie8):

I have posted this question in stackexchange http://math.stackexchange.com/questions/905844/if-dfraca2-b2a2b2-dfrac-sina-b-sinab-then-what-type-o

ganeshie8 (ganeshie8):

Woah ! look at the answer there, we're done !

OpenStudy (anonymous):

right angled triangle ahh

OpenStudy (anonymous):

isosceles right angled triangle

ganeshie8 (ganeshie8):

yes using sine rule, you can replace a/b with sinA/sinB

ganeshie8 (ganeshie8):

a^2/b^2 = tanA cotB sin^2A / sin^2B = tanA cotB

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

(1-cos^2a)/(1-cos^2b)

ganeshie8 (ganeshie8):

don't do that

ganeshie8 (ganeshie8):

a^2/b^2 = tanA cotB sin^2A / sin^2B = tanA cotB sin^2A / sin^2B = (sinAcosB)/(cosAsinB) sinA/sinB = cosB/cosA sinAcosA = sinBcosB sin(2A) = sin(2B) A = B

ganeshie8 (ganeshie8):

so its an `isosceles triangle`

OpenStudy (anonymous):

wait :)

ganeshie8 (ganeshie8):

ha bolo

OpenStudy (anonymous):

yes .thank you so much

ganeshie8 (ganeshie8):

np :)